OpenStudy (anonymous):

Integrate by parts:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x^3 }{(1+x^2)^2 }dx\]

hartnn (hartnn):

the first thing i can think of is to put 1+x^2 as u have u tried it?

OpenStudy (tkhunny):

Really big hint: \[\frac{d}{dx}\frac{1}{1+x^{2}} = ??\]

OpenStudy (anonymous):

hmm let me reformulate my question maybe because I also have a more general one and there was something else I was confused about, though this may be a valid approach. In one solution approach I saw this integral being broke down into \[x^2*\frac{x}{(1+x^2)^2}dx\], then u=x^2 and v'=1+x^2. In this case, the xdx and the squared term over the denominator are not substituted. My question is: can you still do integration by parts in such a case? (Thanks for your input though, I'll try that right now)

OpenStudy (tkhunny):

You'll need an extra 2 in there.

OpenStudy (helder_edwin):

what u proposed looks fine. \[ \large u=x^2\qquad du=2x\,dx \] \[ \large dv=\frac{x}{(1+x^2)^2}\qquad v=-\frac{1}{2}\frac{1}{1+x^2} \]

OpenStudy (helder_edwin):

@nexis . r u still there?

OpenStudy (anonymous):

yes im trying it out right now

OpenStudy (anonymous):

thanks your answer completely cleared this up for me @helder_edwin

OpenStudy (anonymous):

i was confused about what exactly to make dv

OpenStudy (helder_edwin):

u r welcome